Home‎ > ‎Reports‎ > ‎Crew102‎ > ‎

day06


Crew 102 Log Book for March 31, 2011

Photos of the day





Commander's Check-In Report

Franco Carbognani Reporting

Crew Physical Status: improving

Time Departed/Returned from EVA: 18:30-20:00

Brief Narrative of Field Mission Results:

Discovered site extremely abundant in cyanobacteria samples

EVA Data/Interpretations:

Rock Samples analysis ongoing

EVA GPS tracking using crew smart-phones is working well and GPS data are available for download and processing.

Engineering/Hab Maintenance:

Main generator running fine after provisional fixes

Backup (Honey) generator, after some minor fixes, seems back in shape. To be possibly tested in real use with the OK of Mission Support

Water Heater leaking problem manageable.

Propane refill has been done.

Plans for Tomorrow:

Long EVA with ATV for geological sampling.

Short EVA around the Hab dedicated to corrosion inspection and sample collection for optical microscopy analysis

Proceed on analysis of Radio-telescope data after calibration. Verify the possibility to define on-line analysis scripts able to detect energetic solar storms.

Possibly proceeding on bringing back gray water system in operation based on the Mission Support Chief Engineer inputs.

Finalization of provisional Hub Monitoring and Alarming. Interaction with Mission Support is needed in order to understand how to proceed further.

Continue on ongoing psychological and physiological studies.

Support Requested:

Don Lusco for final fixes on power generator.


Commander's Report

Franco Carbognani Reporting

The provisional fix to the power outage problem is holding.

Don did not show up today for a final fix but it seems the main generator is running fine for the moment.

Very productive day on engineering and science side.

Second long EVA has been performed and the Crew has discovered a cyanobacteria "paradise" with plenty of samples for our studies.

The Radio-telescope is satisfactorily up and running and Yuval is collecting plenty of data for off-line analysis.

We are considering the possibility to use the radio-telescope for simulating Solar Storm alerts.

Following Mission Support Chief Engineer detailed instructions, we should have brought back the straining tank into operation with few hours of work in the Green Hab,

We have dismounted and packaged the telescope in order to hand over the mount and the controller to Don Lusco when he will be on-site.

The periodic psychological and  physiological tests we are performing reveled to be both interesting and amusing for the whole Crew.

Hab Structural Corrosion studies are progressing well and samples collection will be started on the coming days.

Mood of the crew is good (even if non cooking days do not help on this side)

Crew member's cold is getting better and so are the stomachaches.

On water issue side we have received reassuring information.

We are just overzealous and curious and we would like to fully understand the "mystery" of the white flakes which can been seen especially if the water is left for a while in the refrigerator.

Tests are ongoing as suggested by Mission Support.


Engineering Report

Franco Carbognani Reporting

Generator/Electricity:

Kitty (Diesel Generator) (Used? Yes/No): Yes

Honey (Gas Generator) (Used? Yes/No): No

Battery Charge Level:

Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):

Kitty provisional repair seems holding. The cooling fan is running smoothly as the whole Kitty engine.

With a bit of electrical maintenance:

1) Cleanup of the generator brushes

2) Check of the breakers and wires

we have been able to bring the backup geverator, Honey, back alive.

The generator now produces the standard AC output Voltages that before was completely absent.

The real test could only be done when trying to power up the Hab.

When Don will be onsite working for the final fixes on Kitty we may want to try this test.


Fuel Status:

Diesel (Full/0.75/0.5/0.25/Empty): 0.70

Propane Tank (Full/0.75/0.5/0.25/Empty): 55%

Gasoline Tank (Full/0.75/0.5/0.25/Empty): 0.70

Notes/Comments:

The Propane has been refilled by the supplier. Thanks!


Oil:

Kitty Oil Quantity (# of Quarts in storage): 4/4

Honey Oil Quantity (# of Quarts in storage):

ATV Oil Quantity (# of Quarts in storage): all 4/4

Car Oil Quantity (# of Quarts in storage): 3/4


Water Status:

Outside Potable Water Tank Level (inches from bottom): 17

Trailer Potable Water Tank Level (inches from bottom): 26

Hab Potable Water Tank Level (inches from bottom): 25

Potable Water Meter Reading: 44261,6

In to GreenHab Meter Reading:

Notes/Comments:

Water heater leaking problem: We have not switched on the heater today


GreenHab:


Condition of plants in Tank 1:

Condition of plants in Tank 2:

Condition of plants in Tank 3:

Greenhab Notes/Comments:

Following Joshua detailed instruction we should have brought back the straining tank into operation.

Waiting for further instructions on how to proceed from there.


Observatory:

Telescope Used? (Yes/No):No

Observatory Notes/Comments:

We have dismounted and packaged the telescope as instructed by Peter Detterline. Boxes are ready into the Observatory in order to be handed over to Don Lusco


Transportation:

Hab Car (NOT RENTAL)

Used (yes/no): no

Oil Checked (yes/no):yes

Tire Status: good

Odometer at end of day: 123761.4

Notes/Comments on Hab Car:


ATV 1 (Opportunity):

Used: y

Oil Checked: n

Fuel Consumed: 1/2 liter

Tire Status: inflated

Comments:


ATV 2 (Spirit):

Used: no

Oil Checked: n

Fuel Consumed: none

Tire Status: inflated

Comments: this ATV does not start


ATV 3 (Viking I):

Used: y

Oil Checked: n

Fuel Consumed: 1/2 liter

Tire Status: good

Comments:


ATV 4 (Viking II):

Used: y

Oil Checked: n

Fuel Consumed: 1/2 liter

Tire Status: good

Comments:


Heating and Ventilation:

Thermostat setting upstairs (Farenheit): 69

Actual temperature upstairs (Farenheit): 75

Thermostat setting downstairs (Farenheit):55

Actual temperature downstairs (Farenheit): 70


Computers/Networking Infrastructure:

Notes/comments on internet/computers?:


General Engineering Notes/Comments:


Science Report (Corrosion Inspection of the Hab)

Angeliki Kapoglou Reporting

According to literature, there is a critical threshold of relative humidity below which common metals do not corrode because there is insufficient moisture to create an electrolyte film on the surface of the metal. For aluminum, on a freshly prepared surface and in unpolluted air, the critical threshold of relative humidity is about 70%. Mattsson and Lindgren  reported that the atmospheric corrosion rates of aluminum 1100 exposed for 10 and 20 years at Phoenix, AZ (desert atmosphere) was 0.000 and 0.076mm/year, respectively. However, research on the corrosion of aluminum in other desert environments is not well documented. Utah Desert features a dry, arid climate. The dry weather is a result of the state's location in the rain shadow of the Sierra Nevada in California. Most of the lowland areas receive less than 12 inches (305 mm) of precipitation annually. More specifically, Hanksville receives 5.1 inches of precipitation annually. Thus, severe atmospheric corrosion was not expected to occur on aluminum in this atmosphere. Nevertheless, aluminum F6061 at MDRS suffered notable deterioration in the Utah Desert atmosphere.

Although there is no any air pollution at the MDRS site, the weather of strong winds and blowing sand is very frequent in this area. Thus, a great quantity of surface desert soil could deposit on the surface of specimens in such a climate. Dust particles have been reported to play an important role in atmospheric corrosion. Thus, in desert environment, atmospheric corrosion of aluminum is mainly affected by dust particles. A second inspection of the outside structure was contacted during another EVA. More corroded parts were identified, again around the doors and windows. More EVAs will follow to try to collect samples for optical microscopy and clean the corrosion products.


EVA 5 Report

Yuval Brodsky Reporting

Time: 1830-2000

Crew Members: Y. Brodsky (EVA CDR), L. Vimercati, K. Manyapu

Site Location: UTM 12S 0520073 4251636 (NAD 27 CONUS)

Transit Mode: ATV

Objectives: Cyanobacteria sample collection

Results: This site was discovered to be extremely abundant in cyanobacteria samples, and will be revisited in the coming days for further sample collection. Approximately 600 grams of rock samples were collected, with cyanobacteria covering ~40% of the total rock surface area.

Lessons Learned: Site discovery was key for the remainder of the mission. It will be beneficial to conduct ATV training/familiarization around MDRS prior to embarking on future ATV EVAs.


Chefs Report

Yuval Brodsky Reporting

1a) Today was

 a non-cooking day

1b) Was today a special day celebrated at the main meal with special food or activity?

N/A

2)  List the foods served at today's main meal (usually the evening meal), giving full names of each.

Reboiled (leftover) Leonardo da Fettucini

Refried (leftover) Santa fe rice

Tortillas

Re-heated (leftover) Chipotle soup

Reheated (leftover) boiled vegetables

3) List any main meal foods not finished by the crew (leftovers stored for later use or discarded)

Boiled vegetables

4)  Recipes for anything more complicated than following package directions exactly, or rehydrating.  

Added various spices to attempt to reinvigorate meal.

5)  Lessons learned 

6) Comments/questions for Kim andzJean, the food study investigators

Crew is feeling unsatisfied with leftovers from non-cooking days - particularly at the lack of texture in the food (ie almost no chewing required). Crew also feels inadequately nourished on non-cooking days.


Journalist Report

Kavya Manyapu Reporting

The last day of the month for crew 102 ended with a long EVA on the ETVs in search of cyanobacteria. Today’s EVA was commanded by Y.Brodsky accompanied by K.Manyapu and L.Vimericati. It was quite an adventure for K.Manyapu riding the ETV for the first time even after her motorcycle experience. The other two crew had fun watching her go up and down the bumpy terrain of Mars. The cap-com of the EVA was A.Kapoglou.  The physiology and human factors experiments continue to take place.  The entire morning was spent by the crew fixing the grey water system by the crew commander F.Carbognani and the back-up generator by Y.brodsky.  The crew has been packing the telescope mount to be sent back to earth as it needs to be repaired.  Today’s food once again turned out to be a disaster!  The crew believes that the food is making them unhealthy; however the crew is still in good health. 

Another week to go for the crew to survive on this red planet and let’s wish them great luck for any unexpected situations that may show up.  

Clear Skies from crew 102